40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 24 18:44:03 2021, Koji, Update, CDS, FB is writing the frames with a year old date Screen_Shot_2021-08-24_at_18.46.24.png
    Reply  Wed Aug 25 17:31:30 2021, Paco, Update, CDS, FB is writing the frames with a year old date TRX_noise_2021-08-25_17-40-55.pngTRX_TRY_power_spectra.pdf
       Reply  Thu Aug 26 10:10:44 2021, Paco, Update, CDS, FB is writing the frames with a year old date Screenshot_from_2021-08-26_10-09-50.pngTRXTRY_Spectra.pdf
    Reply  Wed Aug 25 18:20:21 2021, Jamie, Update, CDS, GPS time on fb1 fixed, dadq writing correct frames again 
Message ID: 16300     Entry time: Thu Aug 26 10:10:44 2021     In reply to: 16298
Author: Paco 
Type: Update 
Category: CDS 
Subject: FB is writing the frames with a year old date 

[paco, ]

We went over the X end to check what was going on with the TRX signal. We spotted the ground terminal coming from the QPD is loosely touching the handle of one of the computers on the rack. When we detached it completely from the rack the noise was gone (attachment 1).

We taped this terminal so it doesn't touch anything accidently. We don't know if this is the best solution since it is probably needs a stable voltage reference. In the Y end those ground terminals are connected to the same point on the rack. The other ground terminals in the X end are just cut.

We also took the PSD of these channels (attachment 2). The noise seem to be gone but TRX is still a bit noisier than TRY. Maybe we should setup a proper ground for the X arm QPD?


We saw that the X end station ALS laser was off. We turned it on and also the crystal oven and reenabled the temperature controller. Green light immidiately appeared. We are now working to restore the ALS lock. After running XARM ASS we were unable to lock the green laser so we went to the XEND and moved the piezo X ALS alignment mirrors until we maximized the transmission in the right mode. We then locked the ALS beams on both arms successfully. It very well could be that the PZT offsets were reset by the power glitch. The XARM ALS still needs some tweaking, its level is ~ 25% of what it was before the power glitch.

Attachment 1: Screenshot_from_2021-08-26_10-09-50.png  85 kB  Uploaded Thu Aug 26 11:18:31 2021  | Hide | Hide all
Screenshot_from_2021-08-26_10-09-50.png
Attachment 2: TRXTRY_Spectra.pdf  17 kB  Uploaded Thu Aug 26 11:24:20 2021  | Hide | Hide all
TRXTRY_Spectra.pdf
ELOG V3.1.3-